Calculate the amount of heat needed to raise the temperature of 2 liter water form 20°C to the building point. The specific heat capacity of water is 4180J/Kg.C°
1
Expert's answer
2019-10-17T11:09:49-0400
"Q=mc(T_2-T_1)"

"Q=(2)(4180)(100-20)=668800\ J=668.8\ kJ"